number of photons emitted in 5 hours by a 50W sodium lamp is( lambda photon = 5893A°

Grade:12

1 Answers

Vaibhav Singh
7 years ago
Since 1 w = 1 J/s. Therefore 50 w = 50 J/s.We know that energy, E = hc/lambda.E = 1242/(5893) ×1.6×10^-19 ×10 . Therefore E = 3.372 ×10^-19 JSince power = E/t.50 = (3.372/5×3600)nn = 2.66×10^24